OpenStudy (anonymous):

Finding the standard matrix for the linear transformation: T:R^2 ---> R^2 reflects a vector about the line y=-x, then projects that vector onto the y axis and then compresses that vector by a factor of 0.5 in the y direction?

OpenStudy (anonymous):

I was also wondering if it's possible to have a different codomain and domain when doing this like T:R^2---> R^3?

OpenStudy (anonymous):

T(1, 0) = (0, -0.5) T(0, 1) = (0, 0)

OpenStudy (anonymous):

\[ \left[ \begin {array}{cc} 0&0\\ -1/2&0\end {array} \right] \]